OpenStudy (anonymous):

write the standard form of the equation of the line passing through the point (-3, 5) and perpendicular to the line -4x - 3y = -6

OpenStudy (anonymous):

First you have to find the slope. So, -3y = 4x - 6 y = -4/3x + 2 Slope of perpendicular line would be 3/4 Equation is: y = 3/4 x + b Substitute (-3, 5) into the equation and you get 5 = 3/4(-3) + b Then find b and you have your equation